Solve for x. 90=x/x+15
Found 2 solutions by jim_thompson5910, stanbon:
Answer by jim_thompson5910(35256)

90=x/(x+15)

90(x+15)=x

90x+1350=x

90x=x-1350

90x-x=-1350

89x=-1350

x = -1350/89

Answer by stanbon(75887)

90 = x/(x+15)
-----
90(x+15) = x
90x + 135 = x
89x = -135
x = -135/89
=====================
Cheers,
Stan H.